OpenStudy (anonymous):

The function H(t) = -16t2 + vt + s shows the height H (t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second.

OpenStudy (unklerhaukus):

Part A: The height; \(H\), of the object, at a time; \(t\), is \[H(t) = -16t^2 + vt + s \] when the object hits the ground, \(H(t=t_f) = 0\)

OpenStudy (unklerhaukus):

so solve: \[-16t_f^2+vt_f+s=0\] for \(t_f\), with \(s=100\), \(v=60\)

OpenStudy (anonymous):

time to max height ( above s not given) -16t^2=100t -16t=100 t=6.25 sec.